OpenStudy (anonymous):

probabillity India play two matches each with West indies andAustralia.In any match, the probabilities of india getting 0,1,and 2 points are "0.45,0.05,and 0.50"respectively.Assuming that the outcomes are independent, the probability of india getting at least 7 points is a)0.0875 b)1/16 c)0.1125 d)none of these

OpenStudy (anonymous):

guy's help me

OpenStudy (anonymous):

@ParthKohli

OpenStudy (anonymous):

help guy's

OpenStudy (owen3):

Lets label the events in which India plays as an opponent. W1, W2, A1, A2

OpenStudy (anonymous):

ok

OpenStudy (owen3):

how many ways can you add up to 7

OpenStudy (anonymous):

means

OpenStudy (anonymous):

if you are talking about to add four no. and getting 7 than i think this is only 1 way, b'cos we can only get two points in one match, so 2+2+2+1=7

OpenStudy (owen3):

That is correct For exactly 8 points we have 1 case. 2 + 2 + 2 + 2 For exactly 7 points we have 4 cases. 2 + 2 + 2 + 1 2 + 2 + 1 + 2 2 + 1 + 2 + 2 1 + 2 + 2 + 2

OpenStudy (owen3):

now replace the cases with the probabilities

OpenStudy (owen3):

P( x = 8) = 0.5*0.5*0.5*0.5 PX = 7) = 4 choose 3 * 0.5^3 * 0.1^1

OpenStudy (anonymous):

but when you calculate for seven points it also count repets that 2 are same so why we count more in 7 points

OpenStudy (owen3):

Am I allowed to give the answer?

OpenStudy (anonymous):

at first you solve and give me answer than we discuss what you say

OpenStudy (owen3):

yes there are 4 ways you can get 2,2,2,1 points

OpenStudy (anonymous):

so what's the answer you get

OpenStudy (owen3):

.1125

OpenStudy (owen3):

i should let you do the calculation

OpenStudy (anonymous):

but the answer was actually in answer pellet was 0.0875

OpenStudy (anonymous):

ah and that is

OpenStudy (owen3):

P ( x is at least 7 ) = P( X >= 7 ) = P(X =7 ) + P(X=8) = (4 C 3) * 0.5^3 * 0.05^1 + 0.5^4 = 4* 0.5^3 * 0.05^1 + 0.5^4 = 0.0875

OpenStudy (anonymous):

ah ok i got it thanks